In a right triangle, the length of the hypotenuse is 20 inches and the legnth of one leg is 15 inches. What is the length of the other leg? This is a multiple choice A. 5 inches B. 5 to the square root of 7 inches C. the square root of 7 inches D. 25 inches? Please help me, thank you very much!
Answer by arunpaul(104)

one leg is A = 20 inch = hypotenuse
2nd leg = B = 15 inch
let the 2nd leg will be x
so x^2+15^2 = 20^2
or x^2 = 400-225 = 175= 25*7
so x = 5 SQRT 7
so the answer is B